OpenStudy (anonymous):

why does log0.01=-2 plz explain

hartnn (hartnn):

can you tell me, 0.01 = \(\large 10^{??}\)

OpenStudy (anonymous):

-2

OpenStudy (anonymous):

oh i see

hartnn (hartnn):

yes, so we have, \(\large \log 10^{-2}\) use the property that, \(\large \log x^m=m \log x\)